A researcher wants to take a random sample of size 40 of a certain population with a mean of 23.5 and a variance of 18.49. What is the approximate standard deviation of sample means?

1
Expert's answer
2022-03-27T07:27:54-0400

"\sigma_{\bar X}=\frac{\sigma}{\sqrt{n}}=\frac{\sqrt{18.49}}{\sqrt{40}}\approx 0.68."